

********************************************************************
UARS WEEKLY STATUS REPORT - 25 June 2004
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
56% of each day's data amounting to about 630 minutes of solar
observing time per day.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2001039. MEPS data files between
2001040 and 2001089 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE was turned off June 19, 2004 as the spacecraft approached the solar
graze period. HALOE gain and boresight calibration tests are being
performed during this yaw cycle with them resuming when HALOE turns on
July 4, 2004 and will continue daily (evaluation during different thermal
conditions) until HALOE is turned off for the low beta spacecraft
operations. HALOE has requested permission to turn on earlier on July 4th
in order to provide measurements coincident with Atmospheric Chemistry
Experiment (ACE)-FTS on the SICAT.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 12 June 2004 (Orbit 69781, GMT Day 164) through
18 June 2004 (Orbit 69886, GMT Day 171).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 164/1302 69789 PEM Z/H TO DAY ONLY
PEM 164/1400 69790 PEM AXIS 1 & 2 ON
The beta angle ranged between the angles of 21.0 degrees to 46.8
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 80 degrees on 27 June 2004 (DOY 180).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.96 to 1.27 degrees C. Additional
battery performance data for each battery on 11 June and 18 June 2004
are provided in Appendix D.
The clock error ranged from 10.4 to -09.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 01
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
PEM = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT was able to recreate the yaw wheel anomaly on the simulator
this week. A single glitch was introduced in the tachometer readings
and the simulator responded exactly like the spacecraft during the
anomaly. This was an important result because it proved that our
theory of a single event upset could certainly have been the cause.
No further occurrences of the glitch have been seen on the spacecraft.
The differential voltage after the yaw rose on batteries 2 and 3, with
2 climbing to 200 on 15 June 2004. Fortunately this level was not
maintained and the differential voltages came back to nominal levels.
The FOT will continue to monitor the differential voltages closely.
MEETINGS
The FOT attended the Eclipse Pre-ship Review on 15 June 2004 to
verify the contents of the new build and determine the shipping data.
The new build will be installed on the prototype next week.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 July 2004 (DOY 199). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 18 June 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
69781 164/0023 04-167 NBTR NOOPS
69789 164/1302 04-169 PEM PEM Z/H TO DAY ONLY
69790 164/1400 04-170 PEM PEM AXIS 1 & 2 ON
69851 168/1456 04-171 OBC/MEMMON CHANGE MEMMON FOR 2 ORBITS
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 18 June 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
SOSTMPRBLFT Y-LO Temperature Bot_Left_Front 69837-69838
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 18 June 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
-None this report period.
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 18 June 2004
11 June 2004 - GMT Day 163 Beta = 17.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-089.6 +5.47/+3.55 +28.0/-14.8 24.5 1.03 19.5
3 +00.0/-123.2 +4.51/+2.60 +26.8/-14.4 24.5 1.02 18.5
18 June 2004 - GMT Day 170 Beta = 46.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-179.2 +4.83/+3.87 +25.2/-14.4 25.4 1.03 22.0
3 +16.8/-112.0 +3.87/+2.60 +24.8/-14.4 25.4 1.02 21.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 18 June 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-032 FOT 69818 166/1148 30832 Generic Late Acquisition
04-032 FOT 69820 166/1436 30832 Generic Late Acquisition
04-033 FOT 69861 169/0702 36148 43 Second Telemetry Dropout
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-032 FOT See List June 04 30832 Generic Late Acquisition
04-033 FOT 69861 169/0702 36148 43 Second Telemetry Dropout